OpenStudy (anonymous):

If f(x) = 2x + 8, find f(9).

sam (.sam.):

f(x) = 2x + 8 f(9) = 2(9) + 8=26

OpenStudy (anonymous):

f(9) = 2(9) + 8 18 + 8 26